The probability that I'll wear a blue shirt tomorrow is 1/4. The probability that I will forget my keys is 4/5. The probability that I'll wear a blue shirt and forget my keys is 1/5. What is the probability that I will wear a blue shirt or forget my keys? (Again - your answer will be a decimal between 0 and 1, give 3 decimal place accuracy)  
Answer #1

P( Blue shirt ) = 1/4

P( Forget Keys) = 4/5

P( Blue shirt and Forget keys) = 1/5

P( Blue shirt OR Forget Keys) = P( Blue shirt) + P( Forget Keys) - P( Blue shirt and Forget keys)

=( 1/4) + ( 4/5) - (1/5)

= 17/20

= 0.850